OpenStudy (anonymous):

slope m of the line that passes through the points (3, –8) and (–7, –4

OpenStudy (tkhunny):

How much do the y-values change? How much do the x-values change?

OpenStudy (anonymous):

m= -4-(-8)/ -7-3=-4/10=-2/5